# A Taxonomy of System-Level Attacks on Deep Learning Models in Autonomous Vehicles

### v1

@@ 6.1.2. RQ2 [Attacked Components]
By examining the taxonomy tree, we observe that the DL models under attack consist of object detection & tracking, steering angle prediction, lane detection, and traffic control models. Most of them support the perception tasks carried out by the AV, by identifying relevant features in the perceived input and manipulating them to cause a mis-prediction. Often, the perceived input is a camera image (ma2024slowtrack, ; yoon2023learning, ; hanfeld2023kidnapping, ; piazzesi2021attack, ; jha2020ml, ), making it possible to reuse the wide range of attacks investigated in previous works on model-level evasion against images, achieved via adversarial noise patches (biggio2018wild, ). This is confirmed by the knowledge required by these attacks, which is white-box at the model level and black-box at the system level (see Figures 3(e), 3(f)). Hence, the problem becomes how to manipulate the simulation environment in which the AV operates in order to introduce the attack patches that are known to trigger a mis-prediction at the model level, assuming that the induced error will propagate inside the system, which is known only as a black box, to eventually cause a system failure.
The systems under attack include Baidu Apollo, CARLA’s driving agent, Tesla, and several other AV systems. However, most attacked systems are executed within a simulator (see Figure 3(c)). There are several practical reasons to prefer simulated AVs over real ones when experimenting with new attacks against AVs, including: (1) simulations are cheaper, faster and not affected by safety constraints; (2) simulations are less constrained and offer higher flexibility in the manipulation of the environment, which could be very expensive or infeasible in the real world; (3) simulations provide more information about the execution, e.g., obtained by monitoring and logging different types of execution/simulation data, which is useful to fine-tune an attack. Such information might be unavailable in the real world.
{mdframed}[backgroundcolor=cyan!15] RQ2 [Attacked Components]: The most attacked components are image-processing neural networks, for which several model-level attacks are readily available in the literature. The problem is then how to manipulate the environment such that the model-level input resembles an adversarial input and how to ensure the propagation of a model-level error to a system-level failure. This seems easier to achieve in simulation than in the real world because a simulator offers higher flexibility, more information, and is generally cheaper, faster, and not potentially dangerous for human beings.
@@ 6.1.3. RQ3 [Threat Models]
The most common assumption is that the attacker has white-box access to a DL model used in an AV system, which is known only as a black-box. Under this assumption, the papers included in our taxonomy explored various threat models:
@@
    Outside Attack: The attacker cannot manipulate the DL model input. Hence, the attacker can only manipulate the environment (von2023deepmaneuver, ; hanfeld2023kidnapping, ; sato2021dirty, ; cao2021invisible, ; jing2021too, ; yang2021robust, ; boloor2020attacking, ; pavlitskaya2020feasibility, ; nassi2020phantom, ) or can interfere with the sensors (yan2022rolling, ; cao2019adversarial, ) to supply an adversarial input to the model.
@@
The above-listed threat models require an increasing amount of capabilities and permissions on the attacker’s side, with the Outside Attack threats being the easiest to mount and the Inside Attack threats the most difficult and challenging. It should be noticed that not all papers are explicit about the underlying threat model they assume, as a result, we often had to infer it from the algorithmic description of the attack. As the risk associated with an attack depends on the damage it can cause, but also on the probability that an attacker can successfully mount the attack, explicit knowledge of the threat model is important to understand and possibly estimate the likelihood of an attack.
{mdframed}[backgroundcolor=cyan!15] RQ3 [Threat Models]: Threat models span from limited capabilities of manipulation of the environment where an AV operates, to the possibility of directly corrupting or replacing input values at inference or even at training time. Such a wide range corresponds to different degrees of attack likelihood and realism. Papers often lack explicit information on the assumed threat model, which makes it difficult for the reader to assess the feasibility of an attack in a given context.
@@ 6.1.4. RQ4 [Consequences]
Most attacks that cause a misprediction of the steering angle result in the vehicle losing its path and sometimes hitting road objects, such as side walls (von2023deepmaneuver, ; piazzesi2021attack, ; sato2021dirty, ; pavlitskaya2020feasibility, ; liu2017trojaning, ). On the other hand, object misdetection often causes the vehicle to fail to see an object or to perceive a fake object as real, leading to unexpected changes in speed or crashes (ma2024slowtrack, ; yoon2023learning, ; zhang2023data, ; piazzesi2021attack, ; cao2021invisible, ; yang2021robust, ; jha2020ml, ; boloor2020attacking, ; nassi2020phantom, ; cao2019adversarial, ). Misclassifications by the model and incorrect decisions can cause the car to ignore traffic signs or collide with other vehicles (yan2022rolling, ; wang2021stop, ).
There are two main chains of attack propagation from the input to a system failure:
@@
    Targeted Propagation: one or more adversarial inputs propagate into the DL model and cause a misprediction, which goes in a specific direction, aimed at producing a specific system-level failure (hanfeld2023kidnapping, ; wang2023does, ; zhang2023data, ; yan2022rolling, ; wang2021stop, ; piazzesi2021attack, ; sato2021dirty, ; cao2021invisible, ; jing2021too, ; jha2020ml, ; boloor2020attacking, ; pavlitskaya2020feasibility, ; nassi2020phantom, ; cao2019adversarial, ; liu2017trojaning, ). For instance, a patch is added to the environment in order to revert the numerical sign of the steering angle prediction, which is supposed to make the car go off its lane (boloor2020attacking, ).
@@
    Untargeted Propagation: one or more adversarial inputs propagate into the DL model and cause an arbitrary misprediction. Propagation to an arbitrary system failure is conjectured and often possible, but it is not targeted explicitly (ma2024slowtrack, ; yoon2023learning, ; von2023deepmaneuver, ; yang2021robust, ; nassi2020phantom, ). For instance, perturbations are added to the input image to cause arbitrary, incorrect bounding boxes. As a result of the model-level misclassification, the self-driving car misbehaves (ma2024slowtrack, ).
Clearly, the targeted propagation chain is more effective as it causes a model-level mis-prediction that is specifically functional to an overall mis-behavior of the whole system. However, such targeted chains require deeper knowledge of the system, which is often unavailable or incomplete. Hence, often papers adopt the untargeted propagation chain and then just observe and evaluate its effectiveness empirically, with no a-priori warranty of success.
{mdframed}[backgroundcolor=cyan!15] RQ4 [Consequences]: There are two main approaches to achieve the system-level consequences of an attack: (1) specific model-level mis-predictions are targeted in specific execution contexts, to ensure the desired consequence at the system level; (2) arbitrary model-level mis-predictions are triggered, and system-level consequences are just observed, but not controlled.
@@ 6.2. Implications for practitioners and researchers
    Transferability of Attacks: Since most attacks have been tested in simulation. While it is safe to assume that these attacks might not be transferable to the real world, Stocco et al. (stocco2022mind, ) found that simulation based failure scenarios allow real-world testing to be anticipated and informed. This is either because the outcomes transfer reliably between simulation and real environments or because the uncertainty profiles in the simulator can help predict the uncertainty levels and outcomes in real-world scenarios.
@@
    Attack Propagation: Wang et al. (wang2023does, ) highlighted that some existing model-level object evasion attacks in AV fail to produce system-level effects. Therefore, it is crucial to analyze the propagation of these attacks, as the effectiveness of a model-level attack on a system depends on the propagation chain it exploits. During our review of related papers, we identified a lack of research addressing why certain attacks on DL models do not escalate to system-level failures. This gap represents a promising area for further investigation in the field.
@@
    Additional Attack Targets: Existing attacks commonly involve creating adversarial objects or fake road lines to mislead the perception components of an AV. These objects are static and do not move. However, Hanfeld et al. (hanfeld2023kidnapping, ) demonstrated an attack using a moving drone, showing that dynamic objects, such as autonomous or human-controlled units, can also be employed by attackers to carry out their attacks.
@@ 8.1. Internal Validity
The query string was crafted with a focus on attacks on autonomous vehicles and refined iteratively based on the number of retrieved papers and the presence of pivotal papers in our results. Running the query string from 2017 to 2024 delivered an initial list of approximately 8,800 papers. Although changing the query and pivot papers would alter this initial number, our subsequent steps — filtering and snowballing — compensate for this effect. So, we expect that any replications would yield results similar to ours.
@@ 9. Conclusion
While model-level adversarial attacks have received huge attention from researchers in the last decades, it is only more recently that researchers considered the possibility of injecting attack vectors into elements of the environment where the system under attack operates, to observe a system-level failure, not just a model-level mis-prediction. The literature on system-level attacks has grown quite rapidly in recent years, going in different directions. Hence, this is the time to organize the existing works into a taxonomy. For this purpose, we retrieved the 19 more relevant papers from the literature and we created a taxonomy to categorize them. Our key findings indicate an unbalanced distribution of papers across taxonomy categories; a prevalence of attacks against image processing components in simulation environments; a variety of threat models, which include black-box attacks and white-box attacks, the latter either at inference or at training time; two main propagation chains, consisting of targeted propagation, aimed at specific system-level failures, and untargeted propagation.
We also discussed a number of implications for practitioners and researchers. In the former case, they include guidelines and suggestions for security testing (e.g., on risk assessment and attack effectiveness). In the latter case, they indicate possible directions for future work, such as: performing attacks on DL components beyond image processing ones; investigating the transferability of attacks from a simulator to the real world; propagation of an attack from model to system; system level security testing for the creation of more secure vehicles, including dynamic entities among the environment elements that contribute to a system-level attack.
@@
The implementations, source code, data, and experimental results are publicly available at https://zenodo.org/records/12806551.
@@ Appendix A Tables with Paper Mapping
[1] Cars SORT(Y5), FairMOT, ByteTrack, BoT-SORT Baidu Apollo in LGSVL simulator
@@
[3] Cars Dave2 An AV vehicle (hopper vehicle) in BeamNG AVs driving past a roadside billboard.
[4] Drones PULP Frontnet
@@
[5] Cars YOLO v2, v3, v5 Baidu Apollo in LGSVL simulator
@@
ation from other cars in Baidu Apollo
@@
[7] Cars Nexar & YOLO v4 Baidu Apollo in LGSVL simulator AVs approaching a traffic light
@@
DRL in decision making
module of SUMO
@@
And intelligent driver model
@@
[9] Cars ResNet-34 Learning by cheating agent in CARLA AVs on a road with pedestrians and other vehicles
[10] Cars OpenPilot’s Automated Lane Centering (ALC)
@@
(A5-L, A5-C) MSF-based AD perception for
in-road obstacle detection in Baidu Apollo v5
Baidu Apollo in LGSVL simulator v5
(multi sensor fusion)
Av on a single lane road
@@
Tesla’s APE lane recognition
(AutoPilot ECU, Electronic Control Unit)
Tesla model S
@@
[13] Cars PointRCNN and PointPillar and PV-RCNN
@@
[14] Cars Kalman Filter in YOLO v3 Baidu Apollo in LGSVL simulator
@@
Two end to end models,
Conditional IL and RL
End to end driving system in CARLA AVs in turns and intersections
[16] Cars DriveNet (an extension of Dave2) Autonomous Driving agent in CARLA AVs driving past a roadside billboard
@@
Captur (equipped with Mobileye 630)
@@
DNN of Lidar-based perception module in
Baidu Apollo
Driving agent in Baidu Apollo AVs driving
[19] Cars End to end NN model of Udacity Simulator Udacity Simulator AVs driving
@@
Ref Attacked Target Attacker’s Capability Attack Type Attacker’s System/Model knowledge Attack/Error Specificity Failure Specificity
@@
Losing detection and subsequently tracking of the
target object

### v2

@@ 7.1.2. RQ2 [Attacked Components]
By examining the taxonomy tree, we observe that the DL models under attack consist of object detection & tracking, steering angle prediction, lane detection, and traffic control models. Most of them support the perception tasks carried out by the AV, by identifying relevant features in the perceived input and manipulating them to cause a mis-prediction. Often, the perceived input is a camera image (ma2024slowtrack, ; yoon2023learning, ; hanfeld2023kidnapping, ; piazzesi2021attack, ; jha2020ml, ), making it possible to reuse the wide range of attacks investigated in previous works on model-level evasion against images, achieved via adversarial noise patches (biggio2018wild, ). This is confirmed by the knowledge required by these attacks, which is white-box at the model level and black-box at the system level (see Figures 3(g), 3(h)). Hence, the problem becomes how to manipulate the simulation environment in which the AV operates in order to introduce the attack patches that are known to trigger a mis-prediction at the model level, assuming that the induced error will propagate inside the system, which is known only as a black box, to eventually cause a system failure.
The systems under attack include Baidu Apollo, CARLA’s driving agent, Tesla, and several other AV systems. However, most attacked systems are executed within a simulator (see Figure 3(e)). There are several practical reasons to prefer simulated AVs over real ones when experimenting with new attacks against AVs, including: (1) simulations are cheaper, faster and not affected by safety constraints; (2) simulations are less constrained and offer higher flexibility in the manipulation of the environment, which could be very expensive or infeasible in the real world; (3) simulations provide more information about the execution, e.g., obtained by monitoring and logging different types of execution/simulation data, which is useful to fine-tune an attack. Such information might be unavailable in the real world.
RQ2 [Attacked Components]: The most attacked components are image-processing neural networks in the perception module, for which several model-level attacks are readily available in the literature. The problem is then how to manipulate the environment such that the model-level input resembles an adversarial input and how to ensure the propagation of a model-level error to a system-level failure. This seems easier to achieve in simulation than in the real world because a simulator offers higher flexibility, more information, and is generally cheaper, faster, and not dangerous for human beings. Finally, almost every attack on MSF systems required a white-box knowledge of the DL model.
@@ 7.1.3. RQ3 [Threat Models]
The most common assumption is that the attacker has white-box access to a DL model used in an AV system, while the system itself is known only as a black-box. Under the assumption of white-box model knowledge and black-box system knowledge, the papers included in our taxonomy explored various threat models:
@@
    Outside Attack: The attacker cannot manipulate the DL model input. Hence, the attacker can only manipulate the environment (liu2024rpau, ; von2023deepmaneuver, ; hanfeld2023kidnapping, ; sato2021dirty, ; cao2021invisible, ; jing2021too, ; yang2021robust, ; boloor2020attacking, ; pavlitskaya2020feasibility, ; nassi2020phantom, ) or can interfere with the sensors (yan2022rolling, ; cao2019adversarial, ) to supply an adversarial input to the model.
@@
Under the assumption of black-box model knowledge, nearly all attacks were Outside Attacks, manipulating the environment, such as the road, with adversarial input (jing2021too, ; boloor2020attacking, ; nassi2020phantom, ). The only exception is an attack that employs a weight injection tool to generate perturbations directly on images (piazzesi2021attack, )
The above-listed threat models require an increasing amount of capabilities and permissions on the attacker’s side, with the Outside Attack threats being the easiest to mount and the Inside Attack threats the most difficult and challenging. It should be noticed that not all papers are explicit about the underlying threat model they assume. As a result, we often had to infer it from the algorithmic description of the attack. As the risk associated with an attack depends not only on the damage it can cause, but also on the probability that an attacker can successfully mount the attack, explicit knowledge of the threat model is important to understand and possibly estimate the likelihood of an attack.
RQ3 [Threat Models]: Threat models span from limited capabilities of manipulation of the environment where an AV operates, to the possibility of directly corrupting or replacing input values at inference or even at training time. Such a wide range corresponds to different degrees of attack likelihood and realism. Papers often lack explicit information on the assumed threat model, which makes it difficult for the reader to assess the feasibility of an attack in a given context. Although modifying the environment might appear easy, most attacks rely on white-box access to the DL model. This makes it less realistic to believe that such attacks could occur in real-world scenarios, unless the model is leaked or the attacker’s surrogate model closely resembles the target model.
@@ 7.1.4. RQ4 [Consequences]
Most attacks that cause a misprediction of the steering angle result in the vehicle losing its path and sometimes hitting road objects, such as side walls (von2023deepmaneuver, ; piazzesi2021attack, ; sato2021dirty, ; pavlitskaya2020feasibility, ; liu2017trojaning, ). On the other hand, object misdetection often causes the vehicle to fail to see an object or to perceive a fake object as real, leading to unexpected changes in speed or crashes (ma2024slowtrack, ; guo2023adv, ; zhang2023data, ; piazzesi2021attack, ; cao2021invisible, ; yang2021robust, ; jha2020ml, ; boloor2020attacking, ; nassi2020phantom, ; cao2019adversarial, ). Misclassifications by the model and incorrect decisions can cause the car to ignore traffic signs or collide with other vehicles (yan2022rolling, ; wang2021stop, ).
There are two main chains of attack consequences from the input to a system failure:
@@
    Targeted Attacks: one or more adversarial inputs are fed into the DL model and cause a misprediction, which goes in a specific direction, aimed at producing a specific system-level failure (liu2024rpau, ; cao2019adversarial, ; zhang2023data, ; yan2022rolling, ; wang2021stop, ; piazzesi2021attack, ; sato2021dirty, ; cao2021invisible, ; jing2021too, ; jha2020ml, ; boloor2020attacking, ; pavlitskaya2020feasibility, ; nassi2020phantom, ; cao2019adversarial, ; liu2017trojaning, ). For instance, a patch is added to the environment in order to revert the numerical sign of the steering angle prediction, which is supposed to make the car go off its lane (boloor2020attacking, ).
@@
    Untargeted Attacks: one or more adversarial inputs are fed into the DL model and cause an arbitrary misprediction. Propagation to an arbitrary system failure is conjectured and often possible, but it is not targeted explicitly (ma2024slowtrack, ; yoon2023learning, ; von2023deepmaneuver, ; yang2021robust, ; nassi2020phantom, ). For instance, perturbations are added to the input image to cause arbitrary, incorrect bounding boxes. As a result of the model-level misclassification, the self-driving car misbehaves (ma2024slowtrack, ).
Clearly, the targeted attack chain is more effective as it causes a model-level mis-prediction that is specifically functional to an overall misbehavior of the whole system. However, such targeted chains require deeper knowledge of the system, which is often unavailable or incomplete. Hence, often papers adopt the untargeted attack chain and then just observe and evaluate its effectiveness empirically, with no warranty of success.
RQ4 [Consequences]: There are two main approaches to achieve the system-level consequences of an attack: (1) specific model-level mis-predictions are targeted in specific execution contexts, to ensure the desired consequence at the system level; (2) arbitrary model-level mis-predictions are triggered, and system-level consequences are just observed, but not controlled.
@@ 7.2. Implications for Practitioners and Researchers
    Vulnerability of Modules and Systems: The perception module is the primary target of most attacks, while other modules, such as planning, localization, and control, are rarely targeted directly. While end-to-end AV systems were expected to be vulnerable to adversarial attacks, as confirmed in the taxonomy, the surprising finding is that MSF systems are also susceptible. However, almost every attack on MSF systems requires white-box knowledge of the DL model and was not designed in a black-box way. We can conclude that, when designing defenses, it is important to consider that the perception module in MSF autonomous vehicles is the most frequently targeted module.
@@
    Attack Propagation: Wang et al. (wang2023does, ) highlighted that some existing model-level object evasion attacks in AV fail to produce system-level effects. Therefore, it is crucial to analyze the propagation of these attacks, as the effectiveness of a model-level attack on a system depends on the propagation chain it exploits. During our review of related papers, we identified a lack of research addressing why certain attacks on DL models do not escalate to system-level failures. This gap represents a promising area for further investigation in the field. Filling this gap to understand why an attack leads to system-level failures can help in identifying both attacks and defenses more efficiently and accurately, potentially leading to more secure AVs.
@@
    Additional Attack Targets: Existing attacks commonly involve creating adversarial objects or fake road lines to mislead the perception components of an AV. These objects are static and do not move, hence, defense techniques should take into account environmental objects and elements more carefully, as they can play a significant role in enabling or triggering attacks. However, Hanfeld et al. (hanfeld2023kidnapping, ) demonstrated an attack using a moving drone, showing that dynamic objects, such as autonomous or human-controlled units, can also be employed by attackers to carry out their attacks. This points to a potential direction for future work, exploring the use of dynamic objects (e.g., moving vehicles or people) as carriers of adversarial patches.
@@ 9.1. Internal Validity
The query string was crafted with a focus on attacks on autonomous vehicles and refined iteratively based on the number of retrieved papers and the presence of pivotal papers in our results. Running the query string from 2017 to 2024 delivered an initial list of approximately 9,700 papers. Although changing the query and pivot papers would alter this initial number, our subsequent steps — filtering and snowballing — compensate for this effect. So, we expect that any replications would yield results similar to ours.
@@ 10. Conclusion
While model-level adversarial attacks have received huge attention from researchers in the last decades, it is only more recently that researchers considered the possibility of injecting attack vectors into elements of the environment where the system under attack operates, to observe a system-level failure, not just a model-level mis-prediction. The literature on system-level attacks has grown quite rapidly in recent years, going in different directions. Hence, this is the time to organize the existing works into a taxonomy. For this purpose, we retrieved 21 more relevant papers from the literature, and created a taxonomy to categorize them. Our key findings indicate an unbalanced distribution of papers across taxonomy categories; a prevalence of attacks against image processing components and the perception module in simulation environments; a variety of threat models, which include black-box attacks and white-box attacks, the latter either at inference or at training time; two main attack chains, consisting of targeted attacks, aimed at specific system-level failures, and untargeted attacks.
We also discussed a number of implications for practitioners and researchers. In the former case, they include guidelines and suggestions for security testing (e.g., on risk assessment and attack effectiveness). In the latter case, they indicate possible directions for future work, such as: performing attacks on DL components beyond image processing ones; designing attacks targeting modules outside perception; addressing the vulnerabilities of both multi and single-sensor fusion systems; investigating the transferability of attacks from a simulator to the real world; propagation of an attack from model to system; system level security testing for the creation of more secure vehicles, including dynamic entities among the environment elements that contribute to a system-level attack.
@@
The implementations, source code, data, and experimental results are publicly available at https://zenodo.org/records/16407709.
@@ Appendix A Tables with Paper Mapping
[1] Cars
SORT(Y5), FairMOT, ByteTrack, BoT-SORT
Baidu Apollo in LGSVL simulator (MSF)
@@
[2] Drones A variation of DroNet (E2E)
Parrot Bebop 2 Drone and
Matlab simulation (SSF)
Drones moving to its destination
[3] Cars
OpenPilot’s Adaptive Cruise
Control ACC system (Perception)
CARLA and Baidu Apollo (MSF) AV driving behind the attacker’s vehicle
@@
[5] Cars Dave2 (E2E)
An AV vehicle (hopper vehicle) in BeamNG
AVs driving past a roadside billboard.
[6] Drones PULP Frontnet (Perception)
@@
[7] Cars YOLO v2, v3, v5 (Perception) Baidu Apollo in LGSVL simulator (MSF)
@@
ation from other cars in Baidu Apollo (MSF)
@@
[9] Cars Nexar & YOLO v4 (Perception) Baidu Apollo in LGSVL simulator (MSF) AVs approaching a traffic light
@@
DRL in decision-making
module of SUMO (Planning)
@@
And intelligent driver model (SSF)
@@
[11] Cars ResNet-34 (E2E) Learning by cheating agent in CARLA (SSF) AVs on a road with pedestrians and other vehicles
[12] Cars
OpenPilot’s Automated Lane Centering
(ALC) (E2E)
@@
(A5-L, A5-C) In-road obstacle detection
Baidu Apollo in LGSVL simulator v5 (MSF) Av on a single lane road
@@
Tesla’s APE lane recognition (AutoPilot
ECU, Electronic Control Unit) (Perception)
Tesla model S (SSF)
@@
[15] Cars
PointRCNN and PointPillar and PV-RCNN
@@
[16] Cars Kalman Filter in YOLO v3 (Perception) Baidu Apollo in LGSVL simulator (MSF)
@@
Conditional Imitation and Reinforcement
learning (E2E) in CARLA
End to end driving system in CARLA (SSF) AVs in turns and intersections
[18] Cars DriveNet (an extension of Dave2) (E2E) Autonomous Driving agent in CARLA (SSF) AVs driving past a roadside billboard
@@
Captur (equipped with Mobileye 630) (MSF)
@@
DNN of the Lidar-based (Perception)
module in Baidu Apollo
Driving agent in Baidu Apollo (MSF) AVs driving
[21] Cars NN model of Udacity Simulator (E2E) Udacity Simulator (SSF) AVs driving
*MSF stands for Multi-Sensor Fusion and SSF stands for Single-Sensor Fusion
@@
Ref Attacked Target Attacker’s Capability Attack Strategy Attacker’s System/Model knowledge Attack/Error Specificity Failure Specificity
@@
[2]
Environment (anything that can
hold a patch horizontally)
Attaching the patch to anything that can
hold it horizontally, facing the drone.
Evasion WM / GS AG / ES Specific
[3] Vehicles and Trucks
Placing the patch on the back of a vehicle
and drive it in front of the AV
Evasion WM / BS AS / ES Specific
@@
Losing detection due to increased latency
and subsequently tracking of the target object
@@
[2]
Misdetection of objects and Misprediction of the
steering wheel angle
Crashing to objects, freezing or going off-route
[3] Misdetection of the vehicle in front Acceleration and crash into the car in front